OpenStudy (anonymous):

need help 5(x + 10) + x

OpenStudy (uri):

5*x + 5*10 + x @vega111

OpenStudy (dmndlife24):

Lets distribute the 5... 5 * x + 5 * 10 + x 5x + 50 + x Add life terms... 6x + 50

OpenStudy (jaceykilla):

Combine like terms: -3x + 3x = 0 -10 + 5x = 10 + 0 -10 + 5x = 10

OpenStudy (anonymous):

ok thank you

OpenStudy (dmndlife24):

lol *like terms ^

OpenStudy (dmndlife24):

No problem

OpenStudy (anonymous):

http://www.wolframalpha.com/ best thing to deal with math

OpenStudy (anonymous):

This is a place where we need to understand math, not look up the answers on a website.